OpenStudy (summersnow8):

For Bismuth (Bi), determine the mass of 139 protons.

OpenStudy (anonymous):

Bismuth usually has 83 protons. Its atomic mass is 208.98 or 209.00 209.00-83 protons=126 neutrons 139 protons+126 neutrons=265 atomic mass of Bi

OpenStudy (summersnow8):

@Shalante the answer is 1.631 x 10^-22 g

OpenStudy (anonymous):

Can you screenshot the question to make sure it is not a series of question relating to this.

OpenStudy (summersnow8):

#34 For Bismuth (Bi), determine Bi → Atomic # = 83 Atomic mass = 208.98 A. The number of atoms in 0.243 g (7.00 x 10^20 atoms) Bi atoms / Avogadro’s # = 0.243 g / 208.98 g Bi atoms = (0.243 g / 208.98 g) * 6.022 x 10^23 Bi atoms = 7.00 x 10^20 B. The mass of 139 protons (1.631 x 10^-22 g)

OpenStudy (ramirocruzo):

Hey....Look its simple, Mass of 1 proton=1.6726219 × 10-27 kilograms, So, Now,\[136\times1.6726219 × 10^{-27} kilograms=2.27\times10^{25}\]

OpenStudy (anonymous):

I did that too but everything seems to fall close at 10^-22 range